OpenStudy (anonymous):

!!!!!!!!!!!!!!!!MEDALS!!!!!!!!!!!!!!!!!!!! In a sample of 60 households, the mean number of cups of milk drank in June was 82. In a much larger study, the standard deviation was determined to be 68. For the following questions, assume the population standard deviation is the same.

OpenStudy (anonymous):

Margin of error=\[1.96\times \frac{ \sigma }{ \sqrt{n} }\]

OpenStudy (anonymous):

1.) Solve for the 95% confidence interval for the mean cups of milk drank in June? Please show all work and explain your steps for full credit

OpenStudy (anonymous):

2.) Explain the difference of how you would solve this if you wanted a 98% confidence interval instead.

OpenStudy (anonymous):

@dan815 @nincompoop